### v1

@@ Acknowledgements
Besides the data behind the figures are available at https://doi.org/10.5281/zenodo.13847488.
@@ Appendix A Catalog structure
| LNU_(U/G/R/I/Z) | float64 | W/Hz | rest-framed luminosity in a given band, |
| rest-framed magnitudes can by derived using | |||
| NUVR,RK,UV,VJ,GR | float64 | AB mag | rest-framed colors in given bands |
| LNU_(U/G/R/I/Z)_ERR | float64 | W/Hz | error of the rest-framed luminosity in a given band |
| (NUVR,RK,UV,VJ,GR)_ERR | float64 | AB mag | error of the rest-framed colors in given bands |
@@
| and/or complex likelihood distribution, e.g.: / | |||
@@
| and/or complex likelihood distribution, e.g.: / | |||
| FLAGOPTICAL | int64 | - | flag to select sources observed with high SNR (SNR ) in optical bands (grz); |
| FLAGOPTICAL = 3(2/1/0): source is observed in 3(2/1/0) band(s) with SNR | |||
| FLAGINFRARED | int64 | - | flag to select sources observed with high SNR (SNR ) in WISE bands (W1, |
@@
| SNR_(G/R/Z/W1-4) | float32 | - | SNR in a given band calculated as: |
@@ Appendix B Stellar mass errors dependence on the quality of the photometry
The errors in the stellar mass and SFR estimates depend on the model assumptions done for the SED fitting and on the spectral coverage included in the SED fits. Standardly, the logarithmic stellar masses are constrained to the dex level (e.g. Conroy 2013; Comparat et al. 2017; Pacifici et al. 2023) depending on the assumptions on the SFH, IMF and other model prescriptions (for our stellar mass estimates the dependence on model assumptions is outlined in Sec. 6). For our stellar mass estimates, there is a clear dependence of the uncertainties on the SNR of the input photometry (see Fig. 17). An accuracy level of dex for galaxies is fulfilled when the g, r, z observations have . In the case of QSO, an accuracy limit of dex is achieved with at least 3 WISE bands measured with (independently on the optical measurements). We note that the errors increase with redshift, implying for QSO at that even with high SNR the errors may reach dex (see the blue tail in the left plot in Fig. 17).
@@ Appendix C Comparison with other catalogs
In this Appendix, we further explore the comparison of the stellar mass obtained with CIGALE for DESI EDR galaxies with the reference catalogs. In Sec. 5 we already discussed the comparison of stellar masses drawn from the COSMOS2020 catalog and AGN and host properties based on the samples overlapping with AGN-COSMOS catalogs. Here, we expand the comparison of the stellar masses with the SDSS(FIREFLY) and SDSS(MPA-JHU) catalogs (Sec. C.1), with the GSWLC (Sec. C.2) and with the DESI catalog presented by Zou et al. 2024 (Sec. C.3).
@@ C.1 SDSS
Up to now, the largest spectroscopic survey, SDSS, provides several VACs including estimates of the stellar properties. In this work, we compare our stellar mass estimates derived with CIGALE with two commonly used SDSS catalogs: i) eBOSS SDSS FIREFLY DR16 SDSS(FIREFLY)141414https://live-sdss4org-dr16.pantheonsite.io/spectro/eboss-firefly-value-added-catalog (Comparat et al. 2017), and ii) MPA-JHU DR8151515https://www.sdss4.org/dr17/spectro/galaxy_mpajhu/ (Kauffmann et al. 2003b; Brinchmann et al. 2004; Tremonti et al. 2004). The SDSS(Firefly) catalog provides stellar properties (age, metallicity, dust reddening, stellar mass, and the SFH) obtained through a comprehensive analysis of the galaxy spectra from SDSS, Baryon Oscillation Spectroscopic Survey (BOSS), and eBOSS. The physical properties were inferred via full spectral fitting with the chi-squared minimization fitting code FIREFLY (Wilkinson et al. 2017). The catalog relies on the stellar population models from Maraston & Strömbäck (2011) utilizing different stellar libraries and IMFs. In this work, we use the stellar masses obtained assuming a Chabrier (2003) IMF and using the MILES libraries (Sánchez-Blázquez et al. 2006; Beifiori et al. 2011; Falcón-Barroso et al. 2011). The SDSS(MPA-JHU) catalog is a VAC of physical properties of SDSS DR8 galaxy derived from fitting their photometry. The catalog provides line measurements, Lick, and other indices as well as the physical properties. The stellar masses and SFRs are derived from the SED fitting (covering ugriz) using the models and methodology outlined in Kauffmann et al. (2003a). To avoid the aperture effect on the stellar mass estimates, these are derived based on the photometry rather than the spectral indices as originally proposed by Kauffmann et al. (2003a). The SED fitting relies on templates generated with Bruzual & Charlot (2003) SSP models and a Kroupa (2001) IMF. The catalog includes the stellar mass measurements corresponding to the median and 2.5%, 16%, 84%, and 97.5% of the PDF.
@@ C.2 GSWLC
In this work, we use the GSWLC X2 catalog (Salim et al. 2018), an updated version of the GSWLC 1 version (Salim et al. 2016). The catalog includes physical properties derived for more than 650,000 galaxies covering 90% of SDSS with the CIGALE SED fitting code. The updated version modifies the energy balance in the SED fitting by using luminosity-dependent IR templates to derive the total IR luminosity from W3 or W4 photometry to derive more accurate SFRs. For SDSS sources classified as AGN based on the BPT diagram (Tremonti et al. 2004), the IR luminosity is corrected before SED fitting (see details in Salim et al. 2018). The GSWLC 2 catalog is based on the Chabrier (2003) and Bruzual & Charlot (2003) SSP models. For the sample of 17,902 DESI galaxies matched with the GSWLC 2 catalog, we find a good agreement between the stellar masses across the full stellar mass range with an offset of the dex towards lower estimates for our measurements (see Fig. 19). Although both catalogs rely on the CIGALE, the dust attenuation and energy balance are different and could explain the offset. Interestingly, the scatter of the stellar mass estimates is correlated with the quality of the WISE photometry (see lower panel in Fig. 19). Sources observed with at least 2 WISE bands with SNR ¿ 3 are characterized by a larger spread. The investigation of this relation is left for future work on the comparison of different stellar mass estimates for DESI galaxies.
@@ C.3 DESI catalogs
The comparison of stellar masses across different VACs built for DESI galaxies will be discussed in future work. We restrict here the comparison to the stellar mass estimates obtained by Zou et al. (2024). Both VACs rely on CIGALE estimates with the same SSP models and IMF, however, they differ in the prescription, e.g: Zou et al. (2024) leave stellar metallicity as a free parameter, while in our catalog it is fixed to a solar value and our catalog accounts for the AGN templates. We compare separately stellar masses for 125,077 star-forming and 9,852 AGN galaxies selected based on the [N II] Baldwin, Phillips, and Terlevich diagram (BPT Baldwin et al. 1981). The discrepancy between stellar mass measurements is higher for AGN than for star-forming galaxies (median difference increases from -0.136 to -0.199 and NMAD from 0.211 to 0.303). The comparison of the stellar mass estimates is shown in Fig. 20.
@@ Appendix D SED fitting modules
As outlined in Sec. 6, we quantify the effect of the model assumptions for the stellar mass estimates on the representative sample of EDR galaxies described in the following App. D.1. The priors of the SED fitting modules that we analyze are summarized in Table 9.
@@
| inclination, i.e. viewing angle | AGNPSY [deg] | 0, 10, 20, 40, 70, 90 |
@@ D.1 Representative sample selection
To analyze the influence of the assumptions made to generate models and the incorporation of the MIR photometry to the SED fitting on the main physical properties of DESI galaxies (stellar masses and SFRs), we select a representative sample of 50,196 sources. The sample consists of seven different galaxy types: BL AGN, NL AGN, composite objects, star-forming galaxies, passive galaxies, retired galaxies, and others. To select the representative sample we restrict the sample of 1,286,124 EDR galaxies (see Sec. 2) to 512,002 sources (GALAXIES and QSOs) spanning out to redshift where H is visible in DESI spectra. The galaxy selection is based on the emission line measurements coming from FastSpecFit v2.0 (Moustakas et al. 2023; Moustakas in prep.). The selection of 1,750 BL AGN is based on the H line requiring that the amplitude of the narrow and broad components have high SNR ( , ) and that the 2-component Gaussian fit is favored over a single fit by (). From the remaining set, NL AGN, composite, and star-forming galaxies are identified based on the [N II] BPT diagram requiring high SNR for the amplitudes of the lines used in the [N II] BPT diagram ( , , ). Following the selection proposed by Kewley et al. (2001) and Kauffmann et al. (2003a), we select 3,900 NL AGN, 8,393 composite, and 101,757 star-forming galaxies. Among the remaining set of galaxies (396,202) we select 47,797 passive galaxies following the criterion based on the equivalent width of the H line:
@@
Sources that did not match any of the above criteria (284,919) are kept as others. Among all the galaxy classes, except for AGN and composite, we randomly select sources in different bins of absolute z magnitude ()-redshift (Z) space. The summary of the composition of the representative sample is given in Table 10.
@@ D.2 The reliability of the fit
The goodness of the fit is quantified by the reduced of the best-fitting model (). In principle, one would expect a good fit with , as indicates issues in fully capturing the data, and points to overfitting. However, since the models are highly non-linear, it is almost impossible to correctly quantify the number of degrees of freedom for each galaxy (Andrae et al. 2010; Chevallard & Charlot 2016), and thus in CIGALE corresponds to the divided by the total number of input fluxes (instead of number of freedom in its statistical definition). Therefore, we follow the commonly used approach of removing the bad fits by sigma-clipping the (e.g. Małek et al. 2018) and keep only those galaxies for which , which corresponds to of the sample. This threshold is also dictated by our visual inspection of the quality of the SED fitting, where we find a degradation of the fitting of the MIR part of the spectrum. We note that in the literature more restrictive cuts () are adapted for the scientific analysis and we encourage the user to use the one more appropriate for their scientific case (e.g. Buat et al. 2021).

### v2

@@ Acknowledgements.
The VAC of physical properties of DESI galaxies is publicly available at https://data.desi.lbl.gov/doc/releases/edr/vac/cigale/. The data behind the figures are available at https://doi.org/10.5281/zenodo.13847488.
@@ Appendix A Catalog structure
| LNU_(U/G/R/I/Z) | float64 | W/Hz | rest-frame luminosity in a given band, |
| rest-frame magnitudes can by derived using | |||
| NUVR,RK,UV,VJ,GR | float64 | AB mag | rest-frame colors in given bands |
| LNU_(U/G/R/I/Z)_ERR | float64 | W/Hz | error of the rest-frame luminosity in a given band |
| (NUVR,RK,UV,VJ,GR)_ERR | float64 | AB mag | error of the rest-frame colors in given bands |
@@
| and/or complex likelihood distribution, e.g., / | |||
@@
| and/or complex likelihood distribution, e.g., / | |||
| FLAGOPTICAL | int64 | - | flag to select sources observed with high S/N (S/N ) in optical bands (grz); |
| FLAGOPTICAL = 3(2/1/0): source is observed in 3(2/1/0) band(s) with S/N | |||
| FLAGINFRARED | int64 | - | flag to select sources observed with high S/N (S/N ) in WISE bands (W1, |
@@
| S/N_(G/R/Z/W1-4) | float32 | - | S/N in a given band calculated as: |
@@ Appendix B Stellar mass errors dependence on the quality of the photometry
The errors in the stellar mass and SFR estimates depend on the model assumptions done for the SED fitting and on the spectral coverage included in the SED fits. Standardly, the logarithmic stellar masses are constrained to the dex level (e.g., Conroy 2013; Comparat et al. 2017; Pacifici et al. 2023) depending on the assumptions on the SFH, IMF and other model prescriptions (for our stellar mass estimates the dependence on model assumptions is outlined in Sect. 6). For our stellar mass estimates, there is a clear dependence of the uncertainties on the S/N of the input photometry (see Fig. 17). An accuracy level of dex for galaxies is fulfilled when the g, r, z observations have . In the case of QSO, an accuracy limit of dex is achieved with at least 3 WISE bands measured with (independently of the optical measurements). We note that the errors increase with redshift, implying for QSO at that even with high S/N the errors may reach dex (see the blue tail in the left plot in Fig. 17).
@@ Appendix C Comparison with other catalogs
In this appendix, we further explore the comparison of the stellar mass obtained with CIGALE for DESI EDR galaxies with the reference catalogs. In Sect. 5, we offer a comparison of stellar masses drawn from the COSMOS2020 catalog and AGN and host properties based on the samples overlapping with AGN-COSMOS catalogs. Here, we expand the comparison of the stellar masses with the SDSS(FIREFLY) and SDSS(MPA-JHU) catalogs (Sect. C.1), with the GSWLC (Sect. C.2) and with the DESI catalog presented by Zou et al. 2024 (Sect. C.3).
@@ C.1 SDSS
To date, the largest spectroscopic survey, SDSS, had provided several VACs including estimates of the stellar properties. In this work, we compare our stellar mass estimates derived with CIGALE with two commonly used SDSS catalogs: i) eBOSS SDSS FIREFLY DR16 SDSS(FIREFLY)191919https://live-sdss4org-dr16.pantheonsite.io/spectro/eboss-firefly-value-added-catalog (Comparat et al. 2017), and ii) MPA-JHU DR8202020https://www.sdss4.org/dr17/spectro/galaxy_mpajhu/ (Kauffmann et al. 2003b; Brinchmann et al. 2004; Tremonti et al. 2004). The SDSS(Firefly) catalog provides stellar properties (age, metallicity, dust reddening, stellar mass, and the SFH) obtained through a comprehensive analysis of the galaxy spectra from SDSS, Baryon Oscillation Spectroscopic Survey (BOSS), and eBOSS. The physical properties were inferred via full spectral fitting with the chi-squared minimization fitting code FIREFLY (Wilkinson et al. 2017). The catalog relies on the stellar population models from Maraston & Strömbäck (2011) utilizing different stellar libraries and IMFs. In this work, we use the stellar masses obtained assuming a Chabrier (2003) IMF and using the MILES libraries (Sánchez-Blázquez et al. 2006; Beifiori et al. 2011; Falcón-Barroso et al. 2011). The SDSS(MPA-JHU) catalog is a VAC of physical properties of SDSS DR8 galaxy derived from fitting their photometry. The catalog provides line measurements, Lick, and other indices as well as the physical properties. The stellar masses and SFRs are derived from the SED fitting (covering ugriz) using the models and methodology outlined in Kauffmann et al. (2003a). To avoid the aperture effect on the stellar mass estimates, these are derived based on the photometry rather than the spectral indices as originally proposed by Kauffmann et al. (2003a). The SED fitting relies on templates generated with Bruzual & Charlot (2003) SSP models and a Kroupa (2001) IMF. The catalog includes the stellar mass measurements corresponding to the median and 2.5%, 16%, 84%, and 97.5% of the PDF.
@@ C.2 GSWLC
In this work, we use the GSWLC X2 catalog (Salim et al. 2018), an updated version of the GSWLC 1 version (Salim et al. 2016). The catalog includes physical properties derived for more than 650,000 galaxies covering 90% of SDSS with the CIGALE SED fitting code. The updated version modifies the energy balance in the SED fitting by using luminosity-dependent IR templates to derive the total IR luminosity from W3 or W4 photometry to derive more accurate SFRs. For SDSS sources classified as AGN based on the BPT diagram (Tremonti et al. 2004), the IR luminosity is corrected before SED fitting (see details in Salim et al. 2018). The GSWLC 2 catalog is based on the Chabrier (2003) and Bruzual & Charlot (2003) SSP models. For the sample of 17,902 DESI galaxies matched with the GSWLC 2 catalog, we find a good agreement between the stellar masses across the full stellar mass range with an offset of the dex towards lower estimates for our measurements (see Fig. 19). Although both catalogs rely on the CIGALE, the dust attenuation and energy balance are different and could explain the offset. Interestingly, the scatter of the stellar mass estimates is correlated with the quality of the WISE photometry (see lower panel in Fig. 19). Sources observed with at least 2 WISE bands with S/N ¿ 3 are characterized by a larger spread. The investigation of this relation is left for future work on the comparison of different stellar mass estimates for DESI galaxies.
@@ C.3 DESI catalogs
The comparison of stellar masses across different VACs built for DESI galaxies will be discussed in future work. We restrict here the comparison to the stellar mass estimates obtained by Zou et al. (2024). Both VACs rely on CIGALE estimates with the same SSP models and IMF, however, they differ in the prescription, e.g: Zou et al. (2024) leave stellar metallicity as a free parameter, while in our catalog it is fixed to a solar value and our catalog accounts for the AGN templates. We compare separately stellar masses for 125,077 star-forming and 9,852 AGN galaxies selected based on the [N II] Baldwin, Phillips, and Terlevich diagram (BPT Baldwin et al. 1981). The discrepancy between stellar mass measurements is higher for AGNs than for star-forming galaxies (median difference increases from -0.136 to -0.199 and NMAD from 0.211 to 0.303). The comparison of the stellar mass estimates is shown in Fig. 20.
@@ Appendix D SED fitting modules
As outlined in Sect. 6, we quantified the effect of the model assumptions for the stellar mass estimates on the representative sample of EDR galaxies described in the following App. D.1. The priors of the SED fitting modules we analyzed are summarized in Table 9.
@@
| inclination, i.e., viewing angle | AGNPSY [deg] | 0, 10, 20, 40, 70, 90 |
@@ D.1 Representative sample selection
To analyze the influence of the assumptions made to generate models and the incorporation of the MIR photometry to the SED fitting on the main physical properties of DESI galaxies (stellar masses and SFRs), we select a representative sample of 50,196 sources. The sample consists of seven different galaxy types: BL AGNs, NL AGNs, composite objects, star-forming galaxies, passive galaxies, retired galaxies, and others. To select the representative sample we restrict the sample of 1,286,124 EDR galaxies (see Sect. 2) to 512,002 sources (GALAXIES and QSOs) spanning out to redshift where H is visible in DESI spectra. The galaxy selection is based on the emission line measurements coming from FastSpecFit v2.0 (Moustakas et al. 2023; Moustakas in prep.). The selection of 1,750 BL AGNs is based on the H line requiring that the amplitude of the narrow and broad components have high S/N ( , ) and that the 2-component Gaussian fit is favored over a single fit by (). From the remaining set, NL AGNs, composite, and star-forming galaxies are identified based on the [N II] BPT diagram requiring high S/N for the amplitudes of the lines used in the [N II] BPT diagram ( , , ). Following the selection proposed by Kewley et al. (2001) and Kauffmann et al. (2003a), we select 3,900 NL AGNs, 8,393 composite, and 101,757 star-forming galaxies. Among the remaining set of galaxies (396,202) we select 47,797 passive galaxies following the criterion based on the equivalent width of the H line:
@@
Sources that did not match any of the above criteria (284,919) are kept as others. Among all the galaxy classes, except for AGNs and composite, we randomly select sources in different bins of absolute z magnitude ()-redshift (Z) space. The summary of the composition of the representative sample is given in Table 10.
@@ D.2 The reliability of the fit
The goodness of the fit is quantified by the reduced of the best-fitting model (). In principle, one would expect a good fit with , as indicates issues in fully capturing the data, and points to overfitting. However, since the models are highly non-linear, it is almost impossible to correctly quantify the number of degrees of freedom for each galaxy (Andrae et al. 2010; Chevallard & Charlot 2016), and thus in CIGALE corresponds to the divided by the total number of input fluxes (instead of number of freedom in its statistical definition). Therefore, we follow the commonly used approach of removing the bad fits by sigma-clipping the (e.g., Małek et al. 2018) and keep only those galaxies for which , which corresponds to of the sample. This threshold is also dictated by our visual inspection of the quality of the SED fitting, where we find a degradation of the fitting of the MIR part of the spectrum. We note that in the literature more restrictive cuts () are adapted for the scientific analysis and we encourage the user to use the one more appropriate for their scientific case (e.g., Buat et al. 2021).